# USP-Gaussian: Unifying Spike-based Image Reconstruction, Pose Correction and Gaussian Splatting

### v1

@@ Abstract
Spike cameras, as an innovative neuromorphic camera that captures scenes with the 0-1 bit stream at 40 kHz, are increasingly employed for the 3D reconstruction task via Neural Radiance Fields (NeRF) or 3D Gaussian Splatting (3DGS). Previous spike-based 3D reconstruction approaches often employ a casecased pipeline: starting with high-quality image reconstruction from spike streams based on established spike-to-image reconstruction algorithms, then progressing to camera pose estimation and 3D reconstruction. However, this cascaded approach suffers from substantial cumulative errors, where quality limitations of initial image reconstructions negatively impact pose estimation, ultimately degrading the fidelity of the 3D reconstruction. To address these issues, we propose a synergistic optimization framework, USP-Gaussian, that unifies spike-based image reconstruction, pose correction, and Gaussian splatting into an end-to-end framework. Leveraging the multi-view consistency afforded by 3DGS and the motion capture capability of the spike camera, our framework enables a joint iterative optimization that seamlessly integrates information between the spike-to-image network and 3DGS. Experiments on synthetic datasets with accurate poses demonstrate that our method surpasses previous approaches by effectively eliminating cascading errors. Moreover, we integrate pose optimization to achieve robust 3D reconstruction in real-world scenarios with inaccurate initial poses, outperforming alternative methods by effectively reducing noise and preserving fine texture details. Our code, data and trained models will be available at https://github.com/chenkang455/USP-Gaussian.
Figure 1: Left. Illustration of our USP-Gaussian framework, where the spike-based image reconstruction network (Recon-Net), camera poses, and 3DGS are collaboratively optimized signified by Refer to caption. Mid. Visual ablation showcasing the performance of Recon-Net and 3DGS with and without (w/ & w/o) the joint optimization, where the input denotes the long-exposure image formulated by accumulating the spike stream over a period as defined in Eq. 6. Right. Training curve comparison for Recon-Net and 3DGS with and without joint optimization.
{NoHyper}* Corresponding authors.
@@ 1 Introduction
Spike cameras [10], inspired by the human retina [44], introduce an innovative paradigm for encoding high-speed dynamic scenes. In contrast to conventional imaging systems [18] that rely on fixed exposure intervals to integrate visual information, spike cameras continuously sense dynamics by recording the spike stream with an impressive firing rate of 40 kHz. Benefiting from the efficient bit representation of spikes, spike cameras can adeptly detect rapid scenes with significantly diminished bandwidth compared to established high-speed cameras such as the Phantom.
Capitalizing on the remarkable capability of the spike camera to capture intricate motion and texture details in high-speed environments, recent research [31, 43, 36, 41, 44, 40] concentrate on its application for high-speed imaging, i.e., establishing the mapping from bit-encoded spike stream to visually perceivable images. Subsequent investigations delve into the estimation of optical flow [38], the enhancement of motion deblur [3, 6, 33] and applications in high-speed object detection [42], semantic segmentation [31], depth estimation [29] and occlusion removal [30].
The Neural Radiance Fields [17] and 3D Gaussian Splatting [11] have emerged as prominent 3D reconstruction technologies in recent years, benefiting from their designed differential training frameworks that map 2D images to 3D representations. 3DGS employs Gaussian primitives for explicit scene representation and the splatting algorithm for projection, thereby achieving faster rendering speed compared to the pixel-by-pixel approach implemented in NeRF. Building on NeRF and 3DGS, the high-speed and continuous capturing potential of the spike camera can be fully unleashed in 3D reconstruction tasks [45, 9, 32].
@@
To this end, SpikeNeRF [45] incorporates a differentiable spike simulator on the outputs of NeRF based on the Spiking Neural Network (SNN), with the loss formulated as the identity between the simulated spike stream and the input spike stream. SpikeGS [32] initiates with a self-supervised Spike-to-Image reconstruction framework built on the Blind Spot Network (BSN) [13], intended to mitigate the performance degradation [4, 5] encountered by supervised learning methods on real-world scenes. With recovered high-quality images, SpikeGS designs a further constraint from the long spike input and integrates it into the 3DGS pipeline. Furthermore, SpikeNVS [7] for rgb-spike multi-modality and Spike-GS [9] for color spike cameras adopt a similar framework, differing only in the input modality. While these works have improved preliminary representation for NeRF/3DGS, they remain constrained within the cascading framework thus inheriting the aforementioned limitation.
To address these challenges, we propose USP-Gaussian, Unifying Spike-based image reconstruction, Pose correction, and Gaussian Splatting within a one-stage optimization framework to mitigate error propagation inherent in three-stage cascading methods. Specifically, USP-Gaussian comprises two synergistic branches as illustrated in Fig. 1, one dedicated to Spike-to-Image reconstruction via Recon-Net and the other to novel view synthesis implemented with 3DGS. To facilitate the self-supervised training of Recon-Net, we design $\mathcal{L}_{\text{recon}}$ based on the constraint between the input spike stream and recovered outputs. In the 3DGS branch, we develop a spike-based 3DGS framework featuring specially designed loss $\mathcal{L}_{\text{gs}}$, which obviates the requirement for pre-recovered high-quality images and accurate poses. While each branch of USP-Gaussian operates independently, we further introduce the joint loss $\mathcal{L}_{\text{joint}}$ derived from the outputs of both 3DGS and Recon-Net to facilitate unified optimization. Visual ablation comparisons of restored images and the training curve are depicted in Fig. 1, demonstrating that the joint optimization of the two branches mutually enhances the recovered textures and the overall training process. We conduct quantitative and qualitative evaluations on the synthetic dataset with accurate pose and the real-world dataset with inaccurate pose estimation. Experiments showcase that our method achieves superior restoration in both datasets, validating the effectiveness of our joint learning framework. In summary, the contributions of our USP-Gaussian are as follows:
@@ 2.1 Spike-based Image Reconstruction
While neuromorphic cameras generate high frame rate bit-stream information, it is imperative to transform them into multi-bit grayscale images that are perceptible to human observers. To achieve this Spike-to-Image target, a plethora of methodologies [31, 36, 41, 44, 40, 4, 5] have been proposed, which can be broadly classified into model-based, supervised, and self-supervised learning paradigms. Model-based methods, rooted in the sampling principle of the spike camera, aim to infer high-fidelity images drawing from the physical spike-image relationship [43], biological retinal models [44], and short-term synaptic plasticity [41]. Confronted with pervasive dark current noise that is challenging to model, learning-based methods are employed to directly deduce high-frame-rate sharp sequences from spike streams with unknown firing patterns. Specifically, Spk2ImgNet [36] designed a hierarchical architecture engineered to capture the intricate temporal dynamics present within spike streams. Zhao et al. [39] introduced an SNN-based image reconstruction framework SSIR, leveraging the bit-stream sequence processing advantage of the SNN. Concurrently, self-supervised learning methodologies based on the BSN [4, 5] have been developed to counteract the performance deterioration inherent in supervised learning methods stemming from the domain disparity between synthetic and real-world datasets. However, the BSN encounters challenges in effectively removing noise, particularly when the noise levels are pronounced, or images contain abundant textural features.
In this paper, we propose a novel self-supervised spike-based image reconstruction framework that fully exploits the multi-view constraints provided by 3D scenes.
@@ 2.2 Novel View Synthesis based on the Neuromorphic Camera
Recent research has extensively employed neuromorphic cameras, including event cameras [8], which encode intensity difference in dynamic scenarios with low-latency and spike cameras [10], for 3D scene reconstruction.
Event Camera. Klenk et al. [12] proposed the E-NeRF based on the event generation mechanism, which can recover sharp 3D scenes under high-speed camera motion. Rudnev et al. [20] further harnessed a color event camera to learn the color representation of scenes. Qi et al. [19] introduced E2NeRF with a multi-modality input, which synthesizes high-quality images of new viewpoints by exploiting the complementary information between the blurry input and the event stream. To fully utilize the motion perception capabilities of event cameras, further research [1, 16] focused on recovering dynamic scenes with rigid deformations from event streams. In addition to NeRF, EvGGS [25] introduces a generalizable Gaussian splatting framework for event cameras designed to reconstruct 3D scenes without retraining. EF-3DGS [14] develops a free-trajectory framework by aligning the RGB input, event streams, and pose. Event3DGS [27] presents an effective solution for rapid ego-motion robotics scenarios.
@@ 3.1 Spike Camera
Figure 3: The working pipeline of our USP-Gaussian. For each camera viewpoint, we initially derive a pose sequence at equal time intervals based on the camera model described in Sec. 4.3. Leveraging the 3DGS rendering pipeline detailed in Sec. 4.2, we subsequently generate a corresponding sharp image sequence for this viewpoint. Concurrently, Recon-Net is employed to reconstruct a sharp sequence from the captured spike stream as outlined in Sec. 4.1. While Recon-Net and 3DGS are trained based on the discrepancy between the re-synthesized blurry image with the input long-exposure image, alignment of their outputs for the same scene is further enforced through joint optimization loss as described in Sec. 4.4. With formulated loss functions, our USP-Gaussian enables the collaborative optimization of Gaussian primitives, camera poses, and the parameters of Recon-Net as formulated in Eq. 5.
@@ 4 Methods
Target Clarification. Given spike streams $\{\mathcal{S}_{v}\}_{v=1}^{V}$ from $V$ different viewpoints with initial inaccurate poses $\{\mathcal{P}_{v}\}_{v=1}^{V}$, the USP-Gaussian is designed to simultaneously optimize the Recon-Net, poses, and Gaussian primitives as shown in Fig. 1, mathematically formulated as:
@@ 4.1 Spike-based Image Reconstruction Network
The Recon-Net $\mathcal{F}$ is designed to facilitate the mapping from the spike input to the sharp image as shown in Fig. 2(b), formulated as $\mathbf{I}(f)=\mathcal{F}(\mathcal{S}_{\mathcal{T}};\theta)$ with $f$ signifying the central moment of the interval $\mathcal{T}$. Prior research [36, 31, 4] employs a fixed and brief window (41 frames in their configuration) of the spike stream as the network input. However, this representation is insufficient for scenarios with low spike firing rates due to the limited information embedded within.
While extending the window size allows the network to encompass more data from the spike stream, it also complicates feature extraction and demands larger computational models for effective processing. To address it, we configure the input of the Recon-Net with a complementary long-short spike stream as shown in Fig. 3.
Given a long spike stream $\mathcal{S}_{\mathcal{T}}$, we extract a short spike stream $\mathcal{S}_{\mathcal{T^{\prime}}}^{t}\subset\mathcal{S}_{\mathcal{T}}$ centered around time $t$, with the network mapping reformulated as $\mathbf{I}(t)=\mathcal{F}(\mathcal{S}_{\mathcal{T}},\mathcal{S}_{\mathcal{T^{\prime}}}^{t},t;\theta)$, where time $t$ is incorporated to guide aligning the short spike within the long spike sequence. Following the event stream pre-processing in [23, 24, 2], we apply voxelization to the long spike stream $\mathcal{S}_{\mathcal{T}}$, reducing its dimensions from $K\times H\times W$ to $\frac{K}{4}\times H\times W$ by accumulating every $4$ frame. The network architecture consists of basic convolutional blocks as illustrated in Fig. 3 and detailed in the supplementary materials.
@@
Motivated by the textural similarity between the blurry image and the long-exposure image, we can employ the motion-reblur loss widely employed in self-supervised motion deblurring frameworks [3, 34, 35, 28] to guide the training of the Recon-Net. Building on the Recon-Net, we reconstruct $M$ frames uniformly across the interval to form the image sequence $\{\mathbf{I}_{\text{rec}}(t_{m})\}_{m=1}^{M}$ and average them to synthesize a blurry image with less noise. Further, we can formulate the loss by comparing the synthesized image to the long-exposure spike input, expressed as:
$\mathbb{L}_{\text{recon}}(M,\mathcal{T})=\mathcal{L}\left(\frac{1}{M}\sum_{m=1}^{M}\mathbf{I}_{\text{rec}}(t_{m}),\mathbf{E}(\mathcal{T})\right),$ (7)
where the loss function $\mathcal{L}$ combines $L_{1}$ loss with a D-SSIM term, defined as $\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}}$, effectively integrating both pixel accuracy and structural similarity.
Nevertheless, the input spike stream $\mathcal{S}_{\mathcal{T}}$ sufficiently captures the information necessary to estimate $\mathbf{E}(\mathcal{T})$ as detailed in Eq. 1, which will lead the Recon-Net to learn a trivial identity mapping. To mitigate it, we fully utilize the continuous representation capability of the spike camera and propose the multiple reblur loss to regularize the reconstructed sequence. Specifically, we extract $N$ sub-intervals from the long exposure $\mathcal{T}_{n}\subset\mathcal{T}$ with the same central moment as shown in Fig. 3 and apply the reblur loss to each sub-interval, formulating the multi-reblur loss as below:
$\mathcal{L}_{\text{recon}}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{L}_{\text{recon}}(M_{n},\mathcal{T}_{n}),$ (8)
@@ 4.4 Joint Optimization
For each pose in the sequence, we select a short spike stream centered around the timestamp $t_{m}$, obtaining the image sequence $\{\mathbf{I}_{\text{rec}}(t_{m})\}_{m=1}^{M}$ based on the Recon-Net in Sec. 4.1 and the image sequence $\{\mathbf{I}_{\text{gs}}(t_{m})\}_{m=1}^{M}$ rendered from the 3DGS in Sec. 4.2 as shown in Fig. 3. While both sequences represent the same scene, the joint optimization loss $\mathcal{L}_{\text{joint}}$ can be formulated as follows:
@@
$\mathcal{L}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{gs}}+\mathcal{L}_{\text{joint}}.$ (14)
$\mathcal{L}_{\text{recon}}$ optimizes Recon-Net parameters, $\mathcal{L}_{\text{gs}}$ refines poses and Gaussian primitives, and $\mathcal{L}_{\text{joint}}$ jointly optimizes all components, with gradient flow shown in Fig. 3.
@@
Figure 5: 3D reconstruction visual comparison of our USP-Gaussian compared with previous methods on the real-world dataset, where the input is the long-exposure image defined in Eq. 6.

### v2

@@ Abstract
Spike camera, as an innovative type of neuromorphic camera that captures scenes with 0-1 bit stream at 40 kHz, is increasingly being employed for the novel view synthesis task building on techniques such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS). Previous spike-based approaches typically follow a three-stage pipeline: I. Spike-to-image reconstruction based on established algorithms. II. Camera poses estimation. III. Novel view synthesis. However, the cascading framework suffers from substantial cumulative errors, i.e., the quality of the initially reconstructed images will impact pose estimation, ultimately limiting the fidelity of the 3D reconstruction. To address this limitation, we propose a synergistic optimization framework USP-Gaussian, which unifies spike-to-image reconstruction, pose correction, and gaussian splatting into an end-to-end pipeline. Leveraging the multi-view consistency afforded by 3DGS and the motion capture capability of the spike camera, our framework enables iterative optimization between the spike-to-image reconstruction network and 3DGS. Experiments on synthetic datasets demonstrate that our method surpasses previous approaches by effectively eliminating cascading errors. Moreover, in real-world scenarios, our method achieves robust 3D reconstruction benefiting from the integration of pose optimization. Our code, data, and trained models are available at https://github.com/chenkang455/USP-Gaussian.
Figure 1: Left. Illustration of our USP-Gaussian framework, where the spike-based image Reconstruction Network (Recon-Net), camera poses, and 3DGS are collaboratively optimized signified by Refer to caption. Mid. Visual ablation showcasing the performance of Recon-Net and 3DGS with and without (w/ & w/o) the joint optimization strategy, with the ablation table depicted in Tab. 4 and the input formulated in Eq. 6. Right. Training curve comparison for Recon-Net and 3DGS with and without joint optimization.
{NoHyper}* Corresponding author.
@@ 1 Introduction
Spike cameras [12], inspired by the human retina [49], introduce an innovative paradigm for encoding high-speed dynamic scenes. In contrast to conventional imaging systems [22] that rely on fixed exposure intervals to integrate visual information, spike cameras continuously sense dynamics by recording the spike stream with an impressive firing rate of 40 kHz. Benefiting from the efficient bit representation of spikes, spike cameras can adeptly detect rapid scenes with significantly diminished bandwidth compared to the high-speed cameras such as the Phantom.
Capitalizing on the remarkable capability of the spike camera to capture intricate motion and texture details in high-speed environments, recent research [36, 48, 41, 45, 49, 44, 5] concentrate on establishing the mapping from bit-encoded spike to visually perceivable images. Subsequent investigations delve into the estimation of optical flow [43], facilitation of motion deblur [3, 8, 38] and applications such as high-speed object detection [46], motion magnification [33], depth estimation [34] and occlusion removal [35].
Neural Radiance Fields [21] and 3D Gaussian Splatting [13] have emerged as prominent novel view synthesis technologies benefiting from their designed differential training frameworks. Among them, 3DGS employs Gaussian primitives for explicit scene representation and the splatting algorithm for projection, thereby achieving faster rendering speed compared to the pixel-by-pixel approach implemented in NeRF. Building on NeRF and 3DGS, the high-speed and continuous capturing potential of the spike camera can be fully unleashed in 3D reconstruction tasks [50, 11, 37].
@@
For example, SpikeGS [37] initiates with a self-supervised spike-to-image reconstruction framework built on the Blind Spot Network (BSN) [15], intended to mitigate the performance degradation [6, 7] encountered by supervised learning methods on real-world scenes. With recovered high-quality images, SpikeGS designs a further constraint from the long spike input and integrates it into the vanilla 3DGS training pipeline. While SpikeGS have improved preliminary representation for 3DGS, it remain constrained within the cascading framework thus inheriting the aforementioned cumulative error problem.
To address these challenges, we propose USP-Gaussian, Unifying Spike-based image reconstruction, Pose correction, and Gaussian Splatting within a one-stage optimization framework to mitigate error propagation inherent in the three-stage cascading pipeline. Specifically, USP-Gaussian comprises two synergistic branches as illustrated in Fig. 1, one dedicated to spike-to-image reconstruction via Recon-Net and the other to novel view synthesis implemented with 3DGS. To facilitate the training of Recon-Net, we design $\mathcal{L}_{\text{rec}}$ based on the constraint between the input spike and recovered outputs. Besides, we supervise the training of 3DGS based on $\mathcal{L}_{\text{gs}}$, which obviates the requirement for pre-recovered high-quality images. While each branch of USP-Gaussian operates independently, we further introduce the joint loss $\mathcal{L}_{\text{joint}}$ derived from the outputs of both two branches to facilitate unified optimization.
Visual ablation comparisons of restored images and the training curve depicted in Fig. 1 demonstrate that the joint optimization of the two branches mutually enhances the recovered textures and the overall training process. We conduct quantitative and qualitative evaluations on the synthetic dataset and the real-world dataset. Experiments showcase that our method achieves superior restoration in both datasets, validating the effectiveness of our joint learning framework. In summary, the contributions of our USP-Gaussian are as follows:
@@ 2.1 Spike-based Image Reconstruction
While neuromorphic cameras generate high frame rate bit-stream information, it is imperative to transform them into multi-bit grayscale images that are perceptible to human observers. To achieve this spike-to-image target, a plethora of methodologies [36, 41, 45, 49, 44, 6, 7, 5] have been proposed, which can be broadly classified into model-based, supervised, and self-supervised learning paradigms. Model-based methods, rooted in the sampling principle of the spike camera, aim to infer high-fidelity images drawing from the physical spike-image relationship [48], biological retinal models [49], and short-term synaptic plasticity [45]. Confronted with pervasive dark current noise that is challenging to model, learning-based methods are employed to directly deduce the sharp image from spike input with unknown firing patterns. Specifically, Spk2ImgNet [41] designed a hierarchical architecture engineered to capture the intricate temporal dynamics present within spike streams. SpikeCLIP [5] utilized the textual description of the captured scene as supervision for training the spike-to-image network. Concurrently, self-supervised learning methodologies based on the BSN [6, 7] are developed to counteract the performance deterioration in supervised learning methods owing to the domain gap between synthetic and real-world datasets.
In this paper, we propose a novel self-supervised spike-based image reconstruction framework leveraging the multi-view constraints provided by 3D scenes.
@@ 2.2 Novel View Synthesis based on the Neuromorphic Camera
Recent research has extensively employed neuromorphic cameras, including event cameras [10, 16, 18, 27, 47], which encode intensity difference in dynamic scenarios with low-latency and spike cameras [12], for 3D scene reconstruction.
Event Camera. Klenk et al. [14] proposed the E-NeRF based on the event generation mechanism, which can recover sharp 3D scenes under high-speed camera motion. Rudnev et al. [23] further harnessed a color event camera to learn the color representation of scenes. [1, 20] focused on recovering dynamic scenes with rigid deformations from event streams. In addition to NeRF, EvGGS [29] introduces a generalizable Gaussian splatting framework for event cameras designed to reconstruct 3D scenes without retraining. EF-3DGS [17] develops a free-trajectory framework by aligning the RGB input, event streams, and pose. Event3DGS [31] presents an effective solution for rapid ego-motion robotics scenarios.
@@ 3.1 Spike Camera
Figure 3: The working pipeline of our USP-Gaussian. For each camera viewpoint, we initially derive a pose sequence at equal time intervals based on the camera model described in Sec. 4.3. Leveraging the 3DGS rendering pipeline detailed in Sec. 4.2, we subsequently generate a corresponding sharp sequence for this viewpoint. Concurrently, Recon-Net is employed to reconstruct the same sequence from the captured spike stream as outlined in Sec. 4.1. While Recon-Net and 3DGS are trained based on the discrepancy between the re-synthesized blurry image with the input long-exposure image, alignment of their outputs for the same scene is further enforced through joint optimization loss as described in Sec. 4.4. With formulated loss functions, our USP-Gaussian enables the collaborative optimization of Gaussian primitives, camera poses, and the parameters of Recon-Net as formulated in Eq. 5.
@@ 4 Methods
Target Clarification. Given the captured spike stream set $\{\mathcal{S}_{v}\}_{v=1}^{V}$ from $V$ different viewpoints with initial inaccurate poses $\{\mathcal{P}_{v}\}_{v=1}^{V}$, the USP-Gaussian aims to optimize the parameters of Recon-Net and Gaussian primitives while outputting refined poses as shown in Fig. 1, formulated as:
@@ 4.1 Spike-based Image Reconstruction Network
The Recon-Net is designed to facilitate the mapping $\mathcal{F}(\cdot)$ from the spike input to the sharp image as shown in Fig. 2(b), formulated as $\mathbf{I}(f)=\mathcal{F}(\mathcal{S}_{\mathcal{T}};\theta)$ with $f$ signifying the central moment of the interval $\mathcal{T}$. Prior research [41, 36, 6] employs a fixed and brief window (41 frames in their configuration) of the spike stream as the network input. However, it is insufficient for scenarios with low spike firing rates due to the limited information embedded within.
While extending the window size allows the network to encompass more features, it complicates the feature extraction process and demands large computational resource. To address it, we configure the input of the Recon-Net with a complementary long-short spike stream as shown in Fig. 3.
Given a long spike stream $\mathcal{S}_{\mathcal{T}}$, we extract a short spike stream $\mathcal{S}_{\mathcal{T^{\prime}}}^{t}\subset\mathcal{S}_{\mathcal{T}}$ centered around time $t$, with the network mapping reformulated as $\mathbf{I}(t)=\mathcal{F}(\mathcal{S}_{\mathcal{T}},\mathcal{S}_{\mathcal{T^{\prime}}}^{t},t;\theta)$, where time $t$ is incorporated to guide the alignment of the short and long spike. Following the event stream pre-processing in [26, 28, 2], we apply voxelization to the long spike stream $\mathcal{S}_{\mathcal{T}}$, reducing its dimensions from $K\times H\times W$ to $\frac{K}{4}\times H\times W$ accumulating every $4$ frames. The network architecture consists of basic convolutional blocks as illustrated in Fig. 3 and detailed in the supplementary materials.
@@
Motivated by the textural similarity between the blurry image and the long-exposure image, we can employ the motion-reblur loss widely employed in self-supervised motion deblurring frameworks [3, 39, 40, 32] to guide the training of the Recon-Net. We initially reconstruct $M$ frames uniformly across the interval to form the image sequence $\{\mathbf{I}_{\text{rec}}(t_{m})\}_{m=1}^{M}$ and further average them to synthesize the reblur image. The loss can be formulated by comparing the reblur image to the long-exposure image, expressed as:
$\mathbb{L}_{\text{rec}}(M,\mathcal{T})=\mathcal{L}\left(\frac{1}{M}\sum_{m=1}^{M}\mathbf{I}_{\text{rec}}(t_{m}),\mathbf{E}(\mathcal{T})\right),$ (7)
where the loss $\mathcal{L}$ [13] combines $L_{1}$ loss with a D-SSIM term, defined as $\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}}$, effectively integrating both pixel accuracy and structural similarity.
Nevertheless, the input spike stream $\mathcal{S}_{\mathcal{T}}$ sufficiently captures the information to estimate $\mathbf{E}(\mathcal{T})$ as detailed in Eq. 6, which will lead the Recon-Net to learn a trivial identity mapping from the spike to the long exposure image. To mitigate it, we propose the multi-reblur loss to regularize the reconstructed sequence. Specifically, we extract $N$ sub-intervals from the long exposure $\mathcal{T}_{n}\subset\mathcal{T}$ with the same central moment and apply the reblur loss to each sub-interval, formulating the multi-reblur loss as below:
$\mathcal{L}_{\text{rec}}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{L}_{\text{rec}}(M_{n},\mathcal{T}_{n}),$ (8)
@@ 4.4 Joint Optimization
For each pose in the sequence, we select a short spike stream centered around the timestamp $t_{m}$, obtaining the image sequence $\{\mathbf{I}_{\text{rec}}(t_{m})\}_{m=1}^{M}$ based on the Recon-Net in Sec. 4.1 and the image sequence $\{\mathbf{I}_{\text{gs}}(t_{m})\}_{m=1}^{M}$ rendered from the 3DGS in Sec. 4.2. While both sequences represent the same scene, the joint optimization loss $\mathcal{L}_{\text{joint}}$ can be formulated as follows:
@@
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gs}}+\mathcal{L}_{\text{joint}}.$ (14)
where $\mathcal{L}_{\text{rec}}$ optimizes Recon-Net parameters, $\mathcal{L}_{\text{gs}}$ refines poses and Gaussian primitives, and $\mathcal{L}_{\text{joint}}$ jointly optimizes all components, with gradient flow shown in Fig. 3.
@@
Figure 5: 3D reconstruction visual comparison of our USP-Gaussian compared with previous methods on the real-world dataset.